OpenStudy (anonymous):

10P6 10P4 i need help to simplify the expression.

OpenStudy (anonymous):

these are my choices. A. 151,200 B. 420 C. 30 D. 5,040

OpenStudy (bibby):

10p6 = 10! ---- (10-6)!

OpenStudy (mathstudent55):

\(_nP_r = \dfrac{n!}{(n - r)!}\)

OpenStudy (anonymous):

huh??? mathstudent i cant see what you put

OpenStudy (bibby):

paste it into here http://www.codecogs.com/latex/eqneditor.php

OpenStudy (mathstudent55):

I wrote it there. Can you see it or do I need to enter it or to send it or something else?

OpenStudy (anonymous):

can you write it on here please

OpenStudy (mathstudent55):

nPr = (n!) / ( (n - r)! ) 10P4 = (10!) / ( 10 - 4)! ) = (10 * 9 * 8 * 7 * 6! ) / ( 6! ) = 10 * 9 * 8 * 7 = 5040

OpenStudy (mathstudent55):

You wrote 10P6 and 10P4. I did 10P4. You can do 10P6 the same way.